Assuming that 1.0g of nahco3 is reacted with excess hcl and completely converted to nacl calculate the mass of nacl obtained


1
Expert's answer
2021-03-16T08:24:36-0400

NaHCO3 + HCl = NaCl + H2CO3

According to the equation, n (NaHCO3) = n (NaCl).

n = m/M

m = n x M

M (NaHCO3) = 84.007 g/mol

M (NaCl) = 58.44 g/mol

n (NaHCO3) = 1 / 84.007 = 0.012 mol

m (NaCl) = 0.012 x 58.44 = 0.7 g